monday 12 4
Download
Skip this Video
Download Presentation
Monday 12.4.

Loading in 2 Seconds...

play fullscreen
1 / 5

Monday 12.4. - PowerPoint PPT Presentation


  • 54 Views
  • Uploaded on

Monday 12.4. 09: 3 0: Physics fill for the day 17:30: Loss of cryo maintain P roblem with the temperature reading of the current lead for RQTL9.67 ( L7). A rtificial jump in temperature (immediately passing from 50 to 100 K ), which lasted for a few seconds.

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about ' Monday 12.4.' - juan


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
monday 12 4
Monday 12.4.
  • 09:30: Physics fill for the day
  • 17:30: Loss of cryo maintain
    • Problem with the temperature reading of the current lead for RQTL9.67 (L7). Artificial jump in temperature (immediately passing from 50 to 100 K), which lasted for a few seconds.
    • Reading shows similar perturbations about every 3 hours with different amplitudes.
    • Problem no yet understood but temporarily ‘solved’ by filtering the reading.
  • 19:00: Request by CMS to have access in cavern for investigating water leak
    • intervention lasted longer than foreseen (over one hour compared to 20min).
    • ALICE also had access in shadow of CMS.
monday 12 41
Monday 12.4.
  • 19:00: Trip of 3 RF modules for Beam 2. OK after reset.
  • 20:50: BIS problem related to RF for Beam2.
    • Latched interlock coming from a short loss of cryo OK.
  • 21:40: SIS problem related to QPSstate of RQX.L8
    • QPS expert confirms QPS of RQX.L8 is OK. Proceed by masking signal for injection and taking shots for beam setup with low intensities.
  • 22:10:XPOC problem with LBDS after dumping pilot beams
    • Missing BCT data.
  • 22:45: Problem related to orbit correction.
  • 23:20: Again XPOCproblem with LBDS after dumping pilot beams.
  • 23:30: Filling in a few injections for bunch length calibration.
  • 00:15: Problems with loading the ramp settings.
    • apparently caused by adjustments during orbit corrections via trims.
monday 12 42
Monday 12.4.
  • Problem related to orbit correction:
      • MCBV.12L4.B1 = 14.54 mrad
      •  MCBV.16L4.B1and MCBV.8L4.B1 = -7.5 mrad
monday 12 43
Monday 12.4.
  • 02:00: Ramp settings OK
  • 02:30: CPOC problem again
  • 03:00: Beams circulating and ready for ramp.
  • 04:00: Ramp finished.
  • 05:00: Stable beams.
slide5
13.4.
  • Plan for the morning:
    • Squeeze in IR1 & IR5.
    • Squeeze in IR8.
    • Establishment of corrected orbit reference followed by re-adjustment of tertiary collimators.
    • Measurement of loss hierarchy.
    • Simulation of asynchronous beam dump.
  • Afternoon:
    • Damper studies at 450 GeV.
  • Night:
    • Physics fill with squeezed optics.
  • Wednesday:
    • Continuation with preparations for higher beam intensities.
ad